# Building an Open AIBOM Standard in the Wild

### v1

@@ Abstract.
Modern software engineering increasingly relies on open, community-driven standards, yet how such standards are created in fast-evolving domains like AI-powered systems remains underexplored. This paper presents a detailed experience report on the development of the AI Bill of Materials (AIBOM) specification, an extension of the ISO/IEC 5962:2021 Software Package Data Exchange (SPDX) software bill of materials (SBOM) standard, which captures AI components such as datasets and iterative training artifacts. Framed through the lens of Action Research (AR), we document a global, multi-stakeholder effort involving over 90 contributors and structured AR cycles. The resulting specification was validated through four complementary approaches: alignment with major regulations and ethical standards (e.g., EU AI Act and IEEE 7000 standards), systematic mapping to six industry use cases, semi-structured practitioner interviews, and an industrial case study. Beyond delivering a validated artefact, our paper documents the process of building the AIBOM specification “in the wild,” and reflects on how it aligns with the AR cycle, and distills lessons that can inform future standardization efforts in the software engineering community.
standardization, AIBOM, SBOM, software bill of materials, Action Research, traceability, transparency, AI governance, AI compliance, trustworthy AI, software supply chain security, SPDX, Software Package Data Exchange, open source community
conference: IEEE/ACM 48th International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP); April 12–18, 2026; Rio De Janeiro, Brazilcopyright: acmcopyrightbooktitle: IEEE/ACM 48th International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP), April 12–18, 2026, Rio De Janeiro, Brazilccs: Computing methodologies Machine learningccs: Software and its engineering Software development techniques
@@ 1. Introduction
Regulators also rely on standards as mechanisms for compliance. In the EU, conformity with the Medical Device Regulation can be shown through harmonised standards such as ISO 13485 (quality management) and IEC 62304 (software lifecycle processes). In the U.S., Software Bill of Material (SBOM) standards like SPDX and CycloneDX are explicitly referenced in federal cybersecurity guidance (of Commerce, 2021; Cybersecurity and Infrastructure Security Agency, 2025) and medical device regulation. These SBOM standards have spurred a plethora of recent work that has examined SBOM adoption, proposed taxonomies, and developed extensions for vulnerability management (Stalnaker et al., 2024; Xia et al., 2024), which underscores their growing importance.
@@
These SBOM standards target traditional software and cannot meaningfully capture AI-specific artefacts such as datasets, trained models, fine-tuned checkpoints, and data pipelines that dominate modern systems (Xia et al., 2023a). To address this limitation, the AI community has introduced transparency frameworks such as model cards (Mitchell et al., 2019), datasheets (Gebru et al., 2021), and factsheets (et al., 2019), as well as machine-readable formats like Croissant (et al., 2024). Yet, these approaches often isolate AI-specific details from the broader software engineering context in which practitioners build and deploy systems. Sculley et al. (Sculley et al., 2015) showed in their seminal work on hidden technical debt that AI components typically form only a small fraction of much larger systems, underscoring the need to document them in situ within the broader software context. Practitioners thus face a critical gap: they lack a unified, standard mechanism to describe AI components as first-class citizens within the software supply chain (Rajbahadur et al., 2021; Hassan et al., 2024).
@@
Our paper presents an experience report on tackling this challenge through the creation of the AI Bill of Materials (AIBOM) specification, an extension to the SPDX SBOM standard. Our SPDX 3.0 extension comprises 36 new fields that treat datasets, models, and their provenance as first-class supply-chain elements to address the trustworthiness challenges of AI systems. Our work makes a distinct contribution by showing how AR can be adapted from its traditional use within single organization (Baskerville, 1999) to guide a global, multi-stakeholder standardization effort “in the wild.” Through this lens, our AIBOM specification (which extends the SBOM standard to represent AI systems) itself emerges as a validated artifact of a disciplined, iterative process. Over 90 contributors participated in diagnose-design-evaluate-reflect cycles, providing a replicable blueprint for developing standards in fast-moving domains in SE, while delivering a practical, community-driven specification.
@@
AIBOM specification has since landed in the SPDX standard, been piloted in industrial settings, and informed regulatory discussions. While the technical specification and schema live in our whitepaper (Bennet et al., 2024), this paper reports our experience in open standardization, detailing our governance, cadence, and pitfalls. We also demonstrate AIBOM specification’s practical utility using four complementary validation methods: (1) regulatory/standards alignment (EU AI Act, US and EU medical-device guidance, IEEE 7000 series) to support compliance practice; (2) systematic mapping to industry use cases (compliance, risk management, supply-chain governance) with coverage analysis; (3) practitioner interviews (n=10) across roles confirming field relevance and surfacing gaps; and (4) a multinational industrial case study showing AIBOM covers ethics/legal checklists, generates portions of model cards, and enables partial automation for third-party AI assets. Together, these results position AIBOM specification as part of an actionable standard and our process as a blueprint for creating standards in fast-moving, multi-stakeholder SE domains.
@@
  • AR at scale. We demonstrate how AR, traditionally used within single organizations (Baskerville, 1999), can be adapted to guide a global, multi-stakeholder standardization effort “in the wild.” Our account documents governance design, release cadence, decision traceability, and iterative cycles that extended SPDX SBOM standard with 36 AI and Dataset specific fields.
  • Validated artifact. We present AIBOM and demonstrate its practicality via four complementary validations.
@@ 2.1. Development and Extension of Software Standards
Standards in SE. Standards are essential in SE as they provide common processes and terminology that minimize ambiguity, enhance interoperability, and ensure the reliable delivery of high-quality systems (Tsui et al., 2022). Broadly, the current standards can be grouped into key categories ranging from lifecycle processes and quality assessment to testing, supply-chain security, and emerging AI-specific considerations, as summarized in Table 1. These standards underpin the structured development, deployment, and governance of modern enterprise software, forming the foundation for consistent engineering practices across diverse domains and organizations.
@@
Standards Development and Extension Pathways. Standards in SE typically evolve through two complementary pathways. The first is the formal route, followed by standards development organizations (SDOs) such as IEEE and ISO/IEC, which emphasize stability, consensus, and global legitimacy. This approach is characterized by well-defined, gated stages including project authorization, working group (WG) formation, draft development, public review, and final balloting (International Organization for Standardization, 2023). These stages typically span 18–48 months for IEEE standards and 24–36 months for ISO/IEC standards (Tsui et al., 2022; IEEE Standards Association, 2023; International Organization for Standardization, 2023). The first version of IEEE POSIX standard, for example, progressed through this process via multiple iterations and ballots before becoming ISO/IEC 9945 (Isaak, 2005).
@@ 2.2. Software Bill of Materials (SBOMs)
From BOM to SBOM. The concept of a Bill of Materials (BOM) originated in manufacturing as a structured inventory of components and sub-assemblies within a product (Jiao et al., 2000). Applying this principle to software, a Software Bill of Materials (SBOM) provides a machine-readable inventory of software components, enabling organisations to understand what software they run, where it originates, and how it can be trusted (Xia et al., 2023b). SBOMs have become central to software supply chain security as modern systems increasingly depend on third-party and open-source components (Stalnaker et al., 2024).
SBOM standards. Three SBOM standards dominate the landscape: SPDX (Foundation, 2025), developed under the Linux Foundation and standardized as ISO/IEC 5962:2021 (International Organization for Standardization and International Electrotechnical Commission, 2021); CycloneDX (OWASP, 2025), introduced by OWASP and standardized as EMCA-424 (Ecma International, 2024); and SWID Tags (ISO/IEC, 2015), maintained by the U.S. National Institute of Standards and Technology (NIST) to provide transparent software identification. While SPDX and CycloneDX are both widely adopted and offer comparable core capabilities, their selection in practice is typically guided by organizational priorities. SPDX is often preferred for license compliance, provenance tracking and governance use cases, while CycloneDX is preferred for vulnerability management, security automation, and CI/CD integration (National Telecommunications and Information Administration (2021), NTIA; Cybersecurity and Infrastructure Security Agency (2023), CISA; OASIS Open, 2023).
Limitations for AI Systems. Despite their value, current SBOMs standards do not capture the full complexity of AI systems (Muirí, 2019). New artifact types, such as datasets and models, introduce deeper provenance, explainability, and compliance requirements. Dataset quality and lineage directly influence performance and fairness (Lu et al., 2022; Rajbahadur et al., 2021); model opacity necessitates metadata on interpretability, hyperparameters, and operational constraints (Barclay et al., 2023); and iterative lifecycles, including pre-training, fine-tuning, and redeployment, create intricate provenance chains and evolving compliance obligations (Hassan et al., 2024). These limitations have led both researchers and practitioners to call for extensions to SBOM standards that address the specific needs of AI systems, laying the foundation for the concept of an AIBOM.
@@ 3.1. Working Group Formation
Motivation and scope. The AI Working Group (WG) was formed in 2021 under the SPDX community to explore how existing SBOM standards could be extended to represent AI systems. Its primary goal was to determine whether SBOMs captured the artifacts and metadata needed for regulatory compliance, license obligations, auditability, and transparency in AI systems.
@@
Outcome 1 By May 2024, shortly after the official release of SPDX 3.0, the WG had held 82 meetings with 92 unique participants, 20 of whom attended more than six sessions. All meetings since April 2022 were recorded and archived publicly for transparency and traceability (contributors, 2024).
@@ 3.2. Defining the AIBOM Specification
Incorporating existing practices. From July 2022, once a comprehensive list had been drafted, members systematically analyzed established documentation artifacts such as Model cards (Mitchell et al., 2019), Datasheets (Gebru et al., 2021), and AI factsheets (et al., 2019) to refine and expand the field set. The goal was not exhaustive mapping, but to identify the most relevant and widely applicable fields recognized as useful and appropriate in the AIBOM context. A detailed mapping of the fields from these sources that were incorporated, excluded and the reasons we did so, is provided in our whitepaper (Bennet et al., 2024).
Balancing adoption and completeness. Throughout this process, the WG prioritised adoption over comprehensiveness by defining a minimal set of required fields—those most likely to exist in real-world projects and satisfy regulatory or auditing needs. Each field had to meet strict inclusion criteria: (1) relevance to AIBOM goals, (2) availability of a representation method, and (3) consensus on its necessity. More ambitious metadata elements were intentionally deferred to future releases. In line with NTIA (National Telecommunications and Information Administration (2021), NTIA) recommendations and emerging best practices, the WG focused on a small, readily adoptable set of required fields. These fields were refined continuously until May 2024 through iterative review and discussion.
@@ 3.3. External Feedback and Public Consultation
Presenting drafts and gathering early feedback. WG members frequently presented draft versions of the AI and Dataset profiles (that form the AIBOM specification along with the rest of the SPDX profiles, please see Section 4 for more details) at several formal venues, including Open Source Summit Europe (2022, 2023), Open Source Summit Japan (2022), and Open Source Summit North America (2023, 2024, 2025) and semi-formal WGs including the OpenChain Licensing WG (Coughlan, 2022), OpenSSF AI WG (Open Source Security Foundation (2025), OpenSSF), and IEEE P7014.1 WG (IEEE P7014.1 Working Group (2024), EEEPG). These events brought together open-source experts, software practitioners, and legal professionals, whose feedback was then discussed in WG meetings and incorporated into subsequent revisions of the specification.
Public release candidates and community review. Starting on 8 May 2023, the WG initiated a year-long public consultation process, releasing two versions of the AIBOM specification: two release candidates (RC1 and RC2) (spd, 2025). The SBOM tooling community was invited to review the specification, model, and profiles and to submit proposed changes as pull requests (PRs) to the public repository. In total, 20 PRs specific to AIBOM were submitted, 16 of which were accepted after detailed discussion in WG meetings.
@@ 4. AIBOM
Figure 2. Overview of SPDX profiles and their extensions for AI and Dataset.:
@@
Building on the methodology described in Section 3, the working group integrated two new profiles:Dataset and AI, into the SPDX 3.0 specification. These extensions expand the SBOM standard beyond traditional software components to capture the artifacts underpinning AI systems. Figure 2 shows how the various SPDX profiles together represent an AIBOM specification (thereby extending the SBOM standard to represent AI systems. We refer to this as either AIBOM standard or AIBOM specification in our paper.
@@
Designed to support regulatory compliance, license attribution, provenance tracking, and responsible development, they remain fully compatible with existing SPDX tooling and workflows. Detailed overview of our AIBOM specification is available in the SPDX 3.0 whitepaper (Bennet et al., 2024) and public repository. [^1] https://github.com/spdx/spdx-3-model As this paper focuses on the process of creating the AIBOM specification rather than the specification itself, we provide only a high-level summary below.
@@
Profiles and extensibility in SPDX. SPDX 3.0 adopts a modular, profile-based architecture that enables domain-specific extensions while preserving interoperability. A profile defines a coherent subset of the model: the Core profile establishes foundational classes and vocabularies, while additional profiles extend these capabilities to represent licensing, build metadata, security, and software composition.
@@
Within this framework, the AIBOM specification for an AI system is realized through the AI and Dataset profiles working alongside existing components (though they can also be used independently to document models or datasets). Rather than existing as standalone artifacts, these profiles extend SPDX in a modular way without fragmenting the broader ecosystem (Figure 2). They build on existing classes and relationships from the Core, Licensing, and Software profiles while introducing domain-specific metadata essential for describing AI artifacts. Together, they allow datasets and models to be represented as first-class supply-chain elements within the same SBOM framework.
@@
AI profile. The AI profile captures AI model-specific components, capturing provenance (identifiers, versions, and licenses), architecture (type, domain, and evaluation metrics), training details (data sources, configurations, and resource usage), and model risks (limitations, safety considerations, and compliance evidence). It provides a comprehensive view of how models are developed, deployed, and maintained.
@@
Dataset profile. The Dataset profile standardizes how datasets are described, capturing provenance (origins, sources, suppliers, and collection processes), descriptive details (size, modality, structure, and preprocessing), intended use (applications and purposes), and risk information (biases, sensitive data, and anonymisation methods). This enables traceability, accountability, and compliance throughout the AI lifecycle.

### v2

@@ Abstract.
Modern software engineering increasingly relies on open, community-driven standards, yet how such standards are created in fast-evolving domains like AI-powered systems remains underexplored. This paper presents a detailed experience report on the development of the AI Bill of Materials (AIBOM) specification, an extension of the ISO/IEC 5962:2021 Software Package Data Exchange (SPDX) software bill of materials (SBOM) standard, which captures AI components such as datasets and iterative training artefacts. Framed through the lens of Action Research (AR), we document a global, multi-stakeholder effort involving over 90 contributors and structured AR cycles. The resulting specification was validated through four complementary approaches: alignment with major regulations and ethical standards (e.g., EU AI Act and IEEE 7000 standards), systematic mapping to six industry use cases, semi-structured practitioner interviews, and an industrial case study. Beyond delivering a validated artefact, our paper documents the process of building the AIBOM specification “in the wild,” and reflects on how it aligns with the AR cycle, and distills lessons that can inform future standardisation efforts in the software engineering community.
standardisation, open source community, Action Research, AIBOM, software bill of materials, SBOM, AI governance, trustworthy AI, software supply chain security, Software Package Data Exchange, SPDX
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIP ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786583.3786921isbn: 979-8-4007-2426-8/2026/04ccs: Computing methodologies Machine learningccs: Software and its engineering Software development techniques
@@ 1. Introduction
Regulators also rely on standards as mechanisms for compliance. In the EU, conformity with the Medical Device Regulation can be shown through harmonised standards such as ISO 13485 (quality management) and IEC 62304 (software lifecycle processes). In the U.S., Software Bill of Material (SBOM) standards like SPDX and CycloneDX are explicitly referenced in federal cybersecurity guidance (of Commerce, 2021; Cybersecurity and Infrastructure Security Agency, 2025a). These SBOM standards have spurred a plethora of recent work that has examined SBOM adoption, proposed taxonomies, and developed extensions for vulnerability management (Stalnaker et al., 2024; Xia et al., 2024), which underscores their growing importance.
@@
These SBOM standards target traditional software and cannot meaningfully capture AI-specific artefacts such as datasets, trained models, fine-tuned checkpoints, and data pipelines that dominate modern systems (Xia et al., 2023b). To address this limitation, the AI community has introduced transparency frameworks such as model cards (Mitchell et al., 2019), datasheets (Gebru et al., 2021), and factsheets (et al., 2019), as well as machine-readable formats like Croissant (et al., 2024). Yet, these approaches often isolate AI-specific details from the broader software engineering context in which practitioners build and deploy systems. Sculley et al. (Sculley et al., 2015) showed in their seminal work on hidden technical debt that AI components typically form only a small fraction of much larger systems, underscoring the need to document them in situ within the broader software context. Practitioners thus face a critical gap: they lack a unified, standard mechanism to describe AI components as first-class citizens within the software supply chain (Rajbahadur et al., 2021; Hassan et al., 2024).
@@
Our paper presents an experience report on tackling this challenge through the creation of the AI Bill of Materials (AIBOM) specification, an extension to the SPDX 3.0 standard. Our extension comprises 36 new fields that treat datasets, models, and their provenance as first-class supply-chain elements to address the trustworthiness challenges. Our work makes a distinct contribution by showing how AR can be adapted from its traditional use within a single organisation (Baskerville, 1999) to guide a global, multi-stakeholder standardisation effort “in the wild.” Through this lens, our AIBOM specification itself emerges as a validated artefact of a disciplined, iterative process. Over 90 contributors participated in diagnose-design-evaluate-reflect cycles, providing a replicable blueprint for developing standards in fast-moving domains in SE, while delivering a practical, community-driven specification.
@@
AIBOM specification has since landed in the SPDX standard, been piloted in industrial settings, and informed regulatory discussions. While the technical specification and schema live in our whitepaper (Bennet et al., 2024), this paper reports our experience in open standardisation, detailing our governance, cadence, and pitfalls. We also demonstrate AIBOM specification’s practical utility using four complementary validation methods: (1) regulatory/standards alignment (EU AI Act, US and EU medical-device guidance, IEEE 7000 series) to support compliance practice; (2) systematic mapping to industry use cases (compliance, risk management, supply-chain governance) with coverage analysis; (3) practitioner interviews (n=10) across roles confirming field relevance and surfacing gaps; and (4) a multinational industrial case study showing AIBOM covers ethics/legal checklists, generates portions of model cards, and enables partial automation for third-party AI assets. Together, these results position AIBOM specification as part of an actionable standard and our process as a blueprint for creating standards in fast-moving, multi-stakeholder SE domains.
@@
  • AR at scale. We demonstrate how AR, traditionally used within single organisations (Baskerville, 1999), can be adapted to guide a global, multi-stakeholder standardisation effort “in the wild.” Our account documents governance design, release cadence, decision traceability, and iterative cycles that extended SPDX SBOM standard with 36 AI and Dataset specific fields.
  • Validated artefact. We present AIBOM and demonstrate its practicality via four complementary validations.
@@ 2.1. Development and Extension of Software Standards
Standards in SE. Standards are vital in SE, providing common processes and terminology that minimise ambiguity, enhance interoperability, and ensure reliable high-quality system delivery (Tsui et al., 2022). Current standards span lifecycle processes, quality assessment, testing, supply-chain security, and emerging AI considerations (Table 1). These frameworks underpin the structured development and governance of modern enterprise software, fostering consistent engineering practices across diverse domains and organisations.
@@
Standards development and extension pathways. Standards in SE typically evolve through two complementary pathways. The first is the formal route, followed by standards development organisations (SDOs) such as IEEE and ISO/IEC, which emphasise stability, consensus, and global legitimacy. This approach is characterised by well-defined, gated stages including project authorisation, working group (WG) formation, draft development, public review, and final balloting (International Organization for Standardization, 2023). These stages typically span 18–48 months for IEEE standards and 24–36 months for ISO/IEC standards (Tsui et al., 2022; IEEE Standards Association, 2023; International Organization for Standardization, 2023). The first version of IEEE POSIX standard, for example, progressed through this process via multiple iterations and ballots before becoming ISO/IEC 9945 (Isaak, 2005).
@@ 2.2. Software Bill of Materials (SBOMs)
From BOM to SBOM. The concept of a Bill of Materials (BOM) originated in manufacturing as a structured inventory of components and sub-assemblies within a product (Jiao et al., 2000). Applying this principle to software, a Software Bill of Materials (SBOM) provides a machine-readable inventory of software components, enabling organisations to understand what software they run, where it originates, and how it can be trusted (Xia et al., 2023a). SBOMs have become central to software supply chain security as modern systems increasingly depend on third-party and open-source components (Stalnaker et al., 2024).
SBOM standards. Three SBOM standards dominate the landscape: SPDX (Foundation et al., 2025b), developed under the Linux Foundation and standardised as ISO/IEC 5962:2021 (International Organization for Standardization and International Electrotechnical Commission, 2021); CycloneDX (OWASP, 2025), introduced by OWASP and standardised as EMCA-424 (Ecma International, 2024); and SWID Tags (ISO/IEC, 2015), maintained by the U.S. National Institute of Standards and Technology (NIST) to provide transparent software identification. While SPDX and CycloneDX are both widely adopted and offer comparable core capabilities, their selection in practice is typically guided by organisational priorities.
Limitations for AI systems. Despite their value, current SBOMs standards do not capture the full complexity of AI systems. New artefact types, such as datasets and models, introduce deeper provenance, explainability, and compliance requirements. Dataset quality and lineage directly influence performance and fairness (Lu et al., 2022; Rajbahadur et al., 2021); model opacity necessitates metadata on interpretability, hyperparameters, and operational constraints (Barclay et al., 2023); and iterative lifecycles, including pre-training, fine-tuning, and redeployment, create intricate provenance chains and evolving compliance obligations (Hassan et al., 2024). These limitations have led both researchers and practitioners to call for extensions to SBOM standards that address the specific needs of AI systems, laying the foundation for the concept of an AIBOM.
@@ 3.1. Working Group Formation
Motivation and scope. The AI Working Group (WG) (Group, 2025) was formed in 2021 under the SPDX community to explore how existing SBOM standards could be extended to represent AI systems. Its primary goal was to determine whether SBOMs captured the artefacts and metadata needed for regulatory compliance, license obligations, auditability, and transparency in AI systems.
@@
Outcome 1 By May 2024, shortly after the official release of SPDX 3.0, the WG had held 82 meetings with 92 unique participants, 20 of whom attended more than six sessions. All meetings since April 2022 were recorded and archived publicly for transparency and traceability (Group, 2024).
@@ 3.2. Defining the AIBOM Specification
Incorporating existing practices. From July 2022, once a comprehensive list had been drafted, members systematically analysed established documentation artefacts such as model cards (Mitchell et al., 2019), datasheets (Gebru et al., 2021), and factsheets (et al., 2019) to refine and expand the field set. The goal was not exhaustive mapping, but to identify the most relevant and widely applicable fields recognised as useful and appropriate in the AIBOM context. A detailed mapping of the fields from these sources that were incorporated, excluded and the reasons we did so, is provided in our whitepaper (Bennet et al., 2024).
Balancing adoption and completeness. Throughout this process, the WG prioritised adoption over comprehensiveness by defining a minimal set of required fields—those most likely to exist in real-world projects and satisfy regulatory or auditing needs. Each field had to meet strict inclusion criteria: (1) relevance to AIBOM goals, (2) availability of a representation method, and (3) consensus on its necessity. More ambitious metadata elements were intentionally deferred to future releases. In line with NTIA (National Telecommunications and Information Administration (NTIA), 2021) recommendations and emerging best practices, the WG focused on a small, readily adoptable set of required fields. These fields were refined continuously until May 2024 through iterative review and discussion.
@@ 3.3. External Feedback and Public Consultation
Presenting drafts and gathering early feedback. WG members frequently presented draft versions of the AI and Dataset profiles (that form the AIBOM specification along with the rest of the SPDX profiles, please see Section 4 for more details) at several formal venues, including Open Source Summit Europe (2022, 2023), Open Source Summit Japan (2022), and Open Source Summit North America (2023, 2024, 2025) and semi-formal WGs including the OpenChain Licensing WG (Coughlan, 2022), OpenSSF AI WG (Open Source Security Foundation, 2025), and IEEE P7014.1 WG (IEEE P7014.1 Working Group (EEEPG), 2024). These events brought together open-source experts, software practitioners, and legal professionals, whose feedback was then discussed in WG meetings and incorporated into subsequent revisions of the specification.
Public release candidates and community review. Starting on 8 May 2023, the WG initiated a year-long public consultation process, releasing two versions of the AIBOM specification: two release candidates (RC1 and RC2) (Foundation et al., 2025a). The SBOM tooling community was invited to review the specification, model, and profiles and to submit proposed changes as pull requests (PRs) to the public repository. In total, 20 PRs specific to AIBOM were submitted, 16 of which were accepted after detailed discussion in WG meetings.
@@ 4. AIBOM
Refer to caption Diagram showing an overview of SPDX profiles and their extensions for AI and Dataset.
Figure 2. Overview of SPDX profiles and their extensions for AI and Dataset.
@@
Building on the methodology described in Section 3, the working group integrated two new profiles: Dataset and AI, into the SPDX 3.0 specification. These extensions expand the SBOM standard beyond traditional software components to capture the artefacts underpinning AI systems. Figure 2 shows how the various SPDX profiles together represent an AIBOM specification (thereby extending the SBOM standard to represent AI systems. In our paper, we refer to this as the AIBOM standard or AIBOM specification.
@@
Designed to support regulatory compliance, licence attribution, provenance tracking, and responsible development, these extensions remain fully compatible with existing SPDX tooling. A detailed overview of the AIBOM specification is available in the SPDX 3.0 white paper (Bennet et al., 2024) and the public repository. [^1] https://github.com/spdx/spdx-3-model As this paper focuses on the specification’s creation process over its technical details, we provide only a high-level summary below.
@@
Profiles and extensibility in SPDX. SPDX 3.0 adopts a profile-based modular architecture that enables domain-specific extensions while preserving interoperability. A profile defines a coherent subset of the model: the Core profile establishes foundational classes and vocabularies, while additional profiles extend these capabilities to represent licencing, build metadata, security, and software composition.
@@
Within this framework, the AIBOM specification for an AI system is realised through the AI and Dataset profiles working alongside existing components (though they can also be used independently to document models or datasets). Rather than exist as standalone artefacts, these profiles extend SPDX in a modular way without fragmenting the broader ecosystem (Figure 2). They build on existing classes and relationships from the Core, Licensing, and Software profiles while introducing domain-specific metadata essential for describing AI artefacts. Together, they allow datasets and models to be represented as first-class supply-chain elements within the same SBOM framework.
@@
AI profile. The AI profile captures AI model-specific components, capturing provenance (identifiers, versions, and licences), architecture (type, domain, and evaluation metrics), training details (data sources, configurations, and resource usage), and model risks (limitations, safety considerations, and compliance evidence). It provides a comprehensive view of how models are developed, deployed, and maintained.
@@
Dataset profile. The Dataset profile standardises how datasets are described, capturing provenance (origins, sources, suppliers, and collection processes), descriptive details (size, modality, structure, and preprocessing), intended use (applications and purposes), and risk information (biases, sensitive data, and anonymisation methods). This enables traceability, accountability, and compliance throughout the AI lifecycle.